OpenStudy (anonymous):

Solve Each Equations: 1) -8n+4(1=5n)= -6n-14 2) -6n-20=-2+4(1-3n)

OpenStudy (noseboy908):

The most important thing to remember with these type of problems is to use PEMDAS. Do you know what that is, and how to use it?

OpenStudy (anonymous):

Yeah, I do

OpenStudy (anonymous):

it just gets confusing with the parenthesis which is the distributive

OpenStudy (noseboy908):

Well, how would you proceed? What do you think?

OpenStudy (anonymous):

well im gonna start with question 1 first

OpenStudy (noseboy908):

Should the "=" in the first equation be a "+"? (the first one?

OpenStudy (noseboy908):

)

OpenStudy (anonymous):

-8n+4(1+5n) = -6n-14

OpenStudy (noseboy908):

Alright, show me what you'd do first.

OpenStudy (anonymous):

you first do the parenthesis

OpenStudy (anonymous):

which is -8+4x1+4x5n=-6n-14

OpenStudy (noseboy908):

If you ever need to multiply any amount of numbers in a algebra problem, remember to never use "x", unless you do it within a set of "( )", ok? It makes it look like you're adding another variable to the problem. Anyway, you're right, so rewrite it after you multiply the 4 both times.

OpenStudy (anonymous):

okay i did it

OpenStudy (noseboy908):

Tell me what you got.

OpenStudy (anonymous):

i didn't solve it yet but i do have a question. Do you keep on solving it with the Pemdas

OpenStudy (noseboy908):

Actually, I was wrong about that, it's not very important here. It's just a matter of isolating n.

OpenStudy (anonymous):

Oh okay

OpenStudy (noseboy908):

So, what'd you get after you multiply the two parts by 4?

OpenStudy (anonymous):

-8n+4 times 1+4 times 5n=-6n-14

OpenStudy (noseboy908):

Yep, though for simplicity, I'd suggest (in the future) not even writing this step down, and just moving on to the next one with it already distributed. (I only had you do it here to show you why it works out how it does.) So, next step?

OpenStudy (anonymous):

Add 8 to the -8n and the -6n

OpenStudy (noseboy908):

What do you mean?

OpenStudy (anonymous):